# CDR: Customizable Density Ratios of Strong-over-weak LLMs for Preference Annotation

### v2

@@ 3.2 Reward Function Customization
Human preferences are multi-dimensional (e.g., safety, trustworthiness, reliability, faithfulness) (Bai et al., 2022; Wang et al., 2024c; Naseem et al., 2024), and an effective reward function should adapt its criteria according to the specific domain requirements. For example, a chatbot explaining corporate vacation policies should emphasize faithfulness to company policy and the accuracy of its responses, rather than focusing on aspects like conversational style or user engagement. However, vanilla log-density ratio reward function provides a single, aggregated reward signal, merging various, potentially conflicting preference aspects. Therefore, it is crucial to define preferences clearly and provide concrete examples to tailor the contrastive reward signal to the specific domain or aspect of human preference.
@@
We introduce CDR, which specifies the domains for each prompt and incorporates instructions and in-context-learning (ICL) examples to define criteria for positive and negative preferences (see figure 4). Each domain has customized instructions and ICL examples, and we ensure diversity by preparing multiple ICL demonstrations, sampling one randomly for each instruction. Formally, for each original user prompt $x$, we inject ICL examples and domain-specific instructions $\operatorname{T}(x)$ to guide the annotation toward relevant preference dimensions. This is equivalent to adapting the reward function into the following form, incorporating $\operatorname{T}(x)$ before applying the log-density ratio for annotation.
$r_{\text{{{CDR}}}}(x,y)=\log\pi_{\text{strong}}(y\mid\operatorname{T}(x),x)-\log\pi_{\text{weak}}(y\mid\operatorname{T}(x),x).$ (8)
@@
To automate annotation, we introduce a domain router that identifies the most relevant domain for each user query, allowing us to apply the appropriate preference criteria for each example in annotation set. For instance, a sensitive query is routed to a Safety expert, while a math or coding query goes to a Math/Code expert. In this paper, we employ the Mixtral 8x7B Instruct v0.1 model (Jiang et al., 2024) with zero-shot prompting to classify prompts into predefined categories (e.g., safety, reasoning, chat) based on a system prompt and task description.
@@
These in-context examples and instructions serve as both demonstrative and descriptive tools to help define and refine the model’s preference criterion. Example templates we used can be found at figure 3 and figure 4. For domains like safety, instructions should include guidelines on how to avoid risky outcomes, while in domains like math, demonstrating the preference criterion through examples may be more effective. These instructions provide high-level guidance by defining overarching principles or definitions that shape the reward function’s preferences during data annotation.
<s><|im_start|>system {system_prompt} <|im_end|> <|im_start|>user You must carefully understand my question and give a relevant, correct, and logical answer. For example: User: {ICL_query_i} Good Assistant: {chosen_response_i} Bad Assistant: {rejected_response_i} Explanation: {explanation_i} User: {user_query} Good Assistant: <|im_end|>
Figure 4: Few-shot Instruction template to guide rewards.
@@ 4.1 Strong-Over-Weak Reward Annotation
To test the effectiveness of the strong-over-weak reward model setup, we design experiments using various models as $\pi_{\text{strong}}$ and $\pi_{\text{weak}}$ across two model families: Mistral and Llama.
@@
We collect off-the-shelf LLMs with varying levels of human alignment. To approximate each model’s degree of alignment to human preferences, we use ArenaHard (Li et al., 2024), a benchmark demonstrated to yield the highest correlation and separability with real human judgments in ChatArena. We then assess the density ratio reward function by pairing distinct model combinations through RewardBench (Lambert et al., 2024). Each sample in RewardBench consists of a human-verified pair: one chosen and one rejected response. The reward function then assigns annotations by comparing the density ratio scores of these two responses. The final score reflects the accuracy of the reward function’s predictions against human-annotated ground truth.
Our experiment includes base models, supervised fine-tuning (SFT) models, as well as models optimized through different preference-tuning algorithms (mostly obtained from  Meng et al. (2024)). These models are respectively labeled as SFT-objective-name in Figure 5 and span techniques such as KTO (Ethayarajh et al., 2024), CPO (Xu et al., 2024), RRHF (Yuan et al., 2023), ORPO (Hong et al., 2024), SLiC-HF (Zhao et al., 2023), IPO (Azar et al., 2023), RDPO (Park et al., 2024), DPO (Rafailov et al., 2023), SimPO (Meng et al., 2024), and PPO (Ouyang et al., 2022). Among them, KTO directly maximizes the utility of generation, where density ratio is used as implied reward. IPO avoids the assumption that pairwise preference can be replaced with point-wise reward. SLiC-HF and RRHF are ranking losses using likelihood or average likelihood margin. ORPO optimizes the odds ratio margin between winning and losing pairs. CPO uses sequence-likelihood margin and SFT loss together. SimPO optimizes length-regularized avg-likelihood margin. PPO is an online RL approach that optimizes policy using external reward. Among all described optimization methods, only RDPO and DPO directly optimize the implicit reward under the BradleyTerry assumption. RRHF, SLiC-HF, CPO, ORPO, and SimPO do not even include a reference model in their objective, implying that their density ratio reward has no connection to DPO implicit reward theory. The diversity of these models enables us to generalize our findings across alignment objectives.
@@ Results
Our findings, visualized in Figure 1, reveal a strong correlation between the accuracy of the reward function in Equation (7) and the strong-over-weak alignment gap. As the alignment gap widens, the reward function achieves stronger results. When the alignment gap is near zero, the signal becomes noisy, with the RewardBench accuracy approximating 50%, indicative of a random guess. Further details are presented in Figure 5, where each row represents a numerator model and each column a denominator model. Each cell displays the reward function’s RewardBench score. The heatmap illustrates that the choice of denominator model substantially influences reward generalization, and selecting weaker denominator models (e.g., Base or SFT) to ensure a sufficient alignment gap, typically leads to much more effective and stable reward functions.
The finding shows considerable flexibility in constructing density ratio reward. For instance, as shown in Figure 1 (left), SFT-RDPO as the numerator performs well with various checkpoints—such as Base, SFT, KTO, RRHF, SLiC-HF, and IPO—as denominators, producing high reward accuracy likely due to these models being less aligned than RDPO. Conversely, using a stronger model as the denominator with SFT-RDPO as the numerator leads to a noticeable drop in reward accuracy. Finally, when Base or SFT models serve as the denominator, nearly any preference-tuned numerator model yields an effective reward function, underscoring that the key to effective reward performance lies in maintaining a meaningful alignment gap rather than requiring DPO or other preference-specific tuning for the numerator model.
@@ 4.2 Customized Strong-Over-Weak Density Ratio
CDR (Customized Density Ratio) uses customized instructions and ICL examples to improve the fidelity and accuracy of vanilla strong-over-weak density ratio. We show that by making customized instructions for different domain, CDR can significantly improve density ratio reward function’s overall and domain-wise reward performance.
@@
We select Nous-Hermes-2-Mistral-7B-DPO  (NousResearch, ) and OpenHermes-2.5-Mistral-7B as the strong-over-weak combination. Extensive evaluations demonstrate the superior performance of the DPO model, which creates a clear separation from the SFT model and positions this pairing among the top-performing density ratio reward functions on RewardBench.
To guide the density ratio toward customized domains, we designed three sets of tailored instructions aimed at improving reward accuracy in the Safety, Code/Math, and ChatHard domains as defined by RewardBench. The Safety instructions address sensitive or high-risk topics, including ethics, harmful behavior, profanity, and legal issues, to encourage safe and responsible responses. The Code/Math instructions focus on coding tasks and mathematical problem-solving, emphasizing logical reasoning, accuracy, and precision. For complex instruction-following tasks, the ChatHard prompt encourages the model to be detail-oriented and demonstrate nuanced understanding of user input. Each instruction set includes domain-specific guidelines and in-context examples (ICLs) that illustrate positive and negative examples within each domain, helping the reward function to generate more precise scores for each query. To optimize domain alignment, an adaptive router based on a zero-shot prompted LLM assigns the appropriate domain instruction set to each example.
@@
Generative reward 53.0 49.5 48.3 52.1 50.0
density ratio (dpo vs. sft) 92.2 60.5 82.4 73.8 77.2
@@
CDR (safety) 88.3 61.8 91.0 87.7 82.5
CDR (code/math) 91.6 60.1 89.9 89.7 83.0
CDR (chat-hard) 89.1 69.7 89.1 85.9 83.5
CDR (adaptive, chat-hard, oracle) 89.1 69.7 91.0 89.7 84.9
CDR (adaptive, oracle) 92.2 60.5 91.0 89.7 83.4
CDR (adaptive, router) 93.9 56.8 91.0 88.0 82.6
Table 1: Performance on Reward Bench across multiple dimensions (Chat, ChatHard, Safety, and Reasoning). The overall score is the average of these four. RM-Mistral-7B is the strongest in-class trained reward model initialized from mistralai/Mistral-7B-Instruct-v0.2. ArmoRM-Llama-3-8B is a SoTA reward model scoring second on RewardBench by time of writing. GPT-4 and Claude-3.5 are proprietary models serving as examples of LLM-as-a-judge reward functions. To construct the density ratio, we can use a DPO model (Nous-Hermes-2-Mistral-7B-DPO), an SFT model ( OpenHermes-2.5-Mistral-7B), or a Base model (Mistral-7B-v0.1). We denote specific pairings in the format (dpo vs. sft), which, for example, indicates the density ratio between DPO and SFT models. CDR (Customized Density Ratio) applies domain-specific instructions (e.g., safety or code/math or chat-hard) when taking density ratio. Adaptive routing configurations include an "oracle" (ideal routing) and a real-world "router" based on a zero-shot prompted LLM.
@@ Results
The results in Table 1 show a clear benefit of employing customized Density Ratio (CDR) approaches across various dimensions. CDR reward function is shown to consistently outperform density ratio reward without domain-customized instructions. CDR reward optimized for safety achieve a Safety score of 91.0, representing a 7.6-point improvement over uninstructed density ratio baselines. This highlights the benefits of safety-specific guidance in enhancing reward function’s safety considerations. Similarly, CDR tailored for code/math achieves a Reasoning score of 89.7, outperforming GPT-4-turbo and Claude-3.5-sonnet, with a substantial 15.9-point gain over baselines. CDR focused on chat-hard scores 69.7 in ChatHard, reflecting improved reward robustness in challenging dialog contexts.
CDR employing an oracle (idealized routing) provides insights into potential performance gains with dynamic routing. Under ideal conditions, CDR can reach an overall score of 84.9, balancing performance across safety, reasoning, and conversational robustness. Adaptive CDR employing a router (real-world routing) shows actual automated final reward performance. Note that the router employ vanilla density ratio for the general chat domain because it scores the highest in Chat, which is most important to downstream preference alignment.
Overall, CDR outperforms standard density ratio baselines by as much as 5.4 points, showing the advantages of adaptively customized reward functions. Generative reward using the same strong model with an identical instruction set performs near random chance. In contrast, CDR that contrasts the strong model versus a weaker model achieves 82.6 overall. The performance is comparable to LLM-as-a-judge reward from GPT-4-turbo and Claude-3.5-sonnet, and surpasses the best-trained reward based on the Mistral-7B model.
@@ 4.3 Alignment with density ratio annotated data
We have demonstrated in previous sections that CDR can achieve strong results in out-of-distribution reward benchmark. In addition, to show the usefulness of CDR in aligning LLMs, we design an on-policy preference alignment experiment similar to setup in Meng et al. (2024); Dong et al. (2024). It is intended to give head-to-head comparisons between proposed log-density ratio reward functions (7) with SoTA reward functions in how well they can preference align a policy. We use Meta-Llama-3-8B-Instruct as the starting model and SimPO (Meng et al., 2024) as the preference optimization algorithm. Our evaluation methods include AlpacaEval2.0, ArenaHard, and MT-Bench, with details in Appendix A.1.
Preference Data Annotation
@@
CDR (safety) 30.0 44.7 2850 39.4 777 8.0
CDR (code/math) 36.0 33.1 1853 30.4 545 8.2
CDR (adaptive) 40.7 46.1 2229 37.4 643 8.0
Table 2: Alignment performance after SimPO training on the Llama-3-Instruct (8B) model. Reward function is used to annotate the online preference dataset, obtained through Best-of-32 sampling. The first row is the performance of the starting model Llama-3-Instruct (8B) model. The second row is the alignment performance of aligning using a SoTA trained reward function. DPO model indicated is NousResearch/Nous-Hermes-2-Mistral-7B-DPO; SFT model is teknium/OpenHermes-2.5-Mistral-7B; Base model is mistralai/Mistral-7B-v0.1. CDR (Customized Density Ratio) applies domain-specific guidance (e.g., safety or code/math) to the vanilla density ratio reward. Adaptive indicates using a routing system to assign instruction from relevant domain for each example.
@@
We focus on two model pairs to define the log-density ratio reward function: (i) SFT model vs. Base model, and (ii) DPO model vs. SFT model. The first model pair (SFT vs. Base) is chosen because neither model has undergone preference tuning, allowing us to test whether a preference reward can be derived based purely on the overall capability improvement after SFT training. The second model pair (DPO vs. SFT) is selected for its highest reward performance, as shown in Table 1. For the prompt-guided reward function, we experiment with various instruction types: no instructions, safety domain instructions, math/coding domain instructions, and customized instructions tailored to the domain of each input prompt.

### v3

@@ 3.2 Reward Function Customization
Human preferences are multi-dimensional (e.g., safety, trustworthiness, reliability, faithfulness) (Bai et al., 2022; Wang et al., 2024d; Naseem et al., 2024), and an effective reward function should adapt its criteria according to the specific domain requirements. For example, a chatbot explaining corporate vacation policies should emphasize faithfulness to company policy and the accuracy of its responses, rather than focusing on aspects like conversational style or user engagement. However, vanilla log-density ratio reward function provides a single, aggregated reward signal, merging various, potentially conflicting preference aspects.
@@
We introduce Dr.SoW, which offers customized preference criterion for annotating samples from different domains through the use of instructions and in-context-learning (ICL) examples. Each domain has its own sets of instructions and ICL examples, and we ensure diversity by preparing multiple ICL demonstrations, sampling one randomly for each instruction. Formally, for each original user prompt $x$, we inject ICL examples and domain-specific instructions $\operatorname{T}(x)$ to guide the annotation toward relevant preference dimensions. This is equivalent to adapting the reward function into the following form, incorporating $\operatorname{T}(x)$ before applying the log-density ratio for annotation.
$r_{\text{{Dr.SoW}}}(x,y)=\log\frac{\pi_{\text{strong}}(y\mid\operatorname{T}(x),x)}{\pi_{\text{weak}}(y\mid\operatorname{T}(x),x)}.$ (3)
@@
To automate annotation, we introduce a domain router that identifies the most relevant domain for each user query. We then apply appropriate preference criteria to each example in the annotation set. For instance, a sensitive query is routed to a Safety expert, while a math or coding query goes to a Math/Code expert. We use the Mixtral 8x7B Instruct v0.1 model (Jiang et al., 2024) with zero-shot prompting to classify prompts into pre-defined categories (e.g., safety, reasoning, chat) based on a system prompt and task description.
@@
We provide a pool of domain-specific in-context examples and instructions, such as those in Figure 8910 (Appendix D.1). They serve as both demonstrative and descriptive tools to help refine the reward model’s preference criterion. Example templates we used can be found in Figure 3. For domains like safety, instructions should include guidelines on how to avoid risky outcomes, while in domains like math, demonstrating the preference criterion through examples may be more effective. These instructions provide high-level guidance by defining overarching principles that shape the reward function’s preferences during data annotation.
If users wish to automatically discover preference criteria for their target domain, we provide an automated pipeline for generating preference instruction prompts. This reduces manual effort in prompt engineering and enhances the accessibility of our approach. Inspired by D’Oosterlinck et al. (2024), our prompt tuning method iteratively constructs the prompt based on an initial prompt and the user-provided evaluation dataset; see details in Appendix D.1. It achieves performance comparable to manually crafted prompts (see Table 7).
@@ 4.1 Strong-Over-Weak Reward Annotation
We collect model pairs, $\pi_{\text{strong}}$ and $\pi_{\text{weak}}$, from two families—Mistral and Llama. These models exhibit distinct levels of human alignment, as measured by ArenaHard (Li et al., 2024), a benchmark demonstrated to yield the highest correlation and separability with real human judgments in ChatArena. We then assess the density ratio reward function of distinct model combinations through RewardBench (Lambert et al., 2024). Each sample in RewardBench consists of a human-verified pair: one chosen and one rejected response. The reward function then assigns annotations by comparing the density ratio scores of these two responses. The final score reflects the accuracy of the reward function’s predictions against human-annotated ground truth. Our experiment includes base models, supervised fine-tuning (SFT) models, as well as models optimized through different preference-tuning algorithms.
@@ Results
Our findings, visualized in Figure 1, reveal a strong correlation between the accuracy of the reward function in Equation (2) and the strong-over-weak alignment gap. As the alignment gap widens, the reward function achieves stronger results. When the alignment gap is near zero, the signal becomes noisy, with the RewardBench accuracy approximating 50%, indicative of a random guess. Further details are presented in Figure 4, where each row represents a numerator model and each column a denominator model. Each cell displays the reward function’s RewardBench score. The heatmap illustrates that the choice of denominator model significantly impacts reward generalization. Selecting weaker denominator models (e.g., Base or SFT) to ensure a sufficient alignment gap typically results in more effective and stable reward functions.
The experiment also shows considerable flexibility in constructing density ratio reward. For instance, as shown in Figure 1 (left), SFT-RDPO as the numerator performs well with various checkpoints—such as Base, SFT, KTO, RRHF, SLiC-HF, and IPO—as denominators, producing high reward accuracy likely due to these models being less aligned than RDPO. Conversely, using a stronger model as the denominator with SFT-RDPO as the numerator leads to a noticeable drop in reward accuracy. Finally, when Base or SFT models serve as the denominator, nearly any preference-tuned numerator model yields an effective reward function, underscoring that the key to effective reward performance lies in maintaining a meaningful alignment gap rather than requiring DPO or other preference-specific tuning for the numerator model.
@@ 4.2 Customized Strong-Over-Weak Density Ratio
Dr.SoW proposes to use customized instructions and in-context learning (ICL) examples to enhance control and accuracy over the vanilla strong-over-weak density ratio. We examine the effect of prompt-based customization in following experiments.
@@
We select Nous-Hermes-2-Mistral-7B-DPO  (NousResearch, ) and OpenHermes-2.5-Mistral-7B as the model pair in Dr.SoW. To tailor vanilla density ratio to specific domains, we develop three customized instruction sets to enhance reward accuracy in Safety, Code/Math, and ChatHard domains. The Safety set focuses on sensitive or high-risk topics like ethics, harmful behavior, profanity, and legal issues, promoting safe and responsible responses. The Code/Math set targets coding tasks and mathematical problem-solving, prioritizing logical reasoning, accuracy, and precision. The ChatHard set emphasizes detailed, nuanced understanding for complex instruction-following tasks. Each set includes domain-specific guidelines and in-context examples (ICLs) showcasing positive and negative cases, enabling the reward function to produce more precise scores. An adaptive router, powered by a zero-shot prompted LLM, assigns the most relevant instruction set to each sample, improving domain adaptability.
@@
DPO model-as-a-judge 53.0 49.5 48.3 52.1 50.0
@@
vanilla density ratio 92.2 60.5 82.4 73.8 77.2
Dr.SoW (safety) 88.3 61.8 91.0 87.7 82.5
Dr.SoW (code/math) 91.6 60.1 89.9 89.7 83.0
Dr.SoW (chat-hard) 89.1 69.7 89.1 85.9 83.5
Dr.SoW (adaptive, chat-hard, oracle) 89.1 69.7 91.0 89.7 84.9
Dr.SoW (adaptive, oracle) 92.2 60.5 91.0 89.7 83.4
Dr.SoW (adaptive, router) 93.9 56.8 91.0 88.0 82.6
Table 1: Performance on Reward Bench across multiple dimensions (Chat, ChatHard, Safety, and Reasoning). The overall score is the average of these four. RM-Mistral-7B is the strongest in-class trained reward model initialized from mistralai/Mistral-7B-Instruct-v0.2. ArmoRM-Llama-3-8B is a SoTA reward model scoring second on RewardBench by time of writing. GPT-4 and Claude-3.5 are proprietary models serving as examples of LLM-as-a-judge reward functions. To construct the density ratio, we can use a DPO model (Nous-Hermes-2-Mistral-7B-DPO), an SFT model ( OpenHermes-2.5-Mistral-7B), or a Base model (Mistral-7B-v0.1). We denote specific pairings in the format (dpo vs. sft), which, for example, indicates the density ratio between DPO and SFT models. Dr.SoW applies domain-specific instructions (e.g., safety or code/math or chat-hard) when taking density ratio. Adaptive routing configurations include an “oracle” (ideal routing) and a real-world “router” based on a zero-shot prompted LLM.
@@ Results
The results in Table 1 show a clear benefit of employing Dr.SoW approaches across various dimensions. Dr.SoW reward function is shown to consistently outperform vanilla density ratio without domain-customized instructions. Dr.SoW reward optimized for safety achieve a Safety score of 91.0, representing a 7.6-point improvement over uninstructed density ratio baselines. This highlights the benefits of safety-specific guidance in enhancing reward function’s safety considerations. Similarly, Dr.SoW tailored for code/math achieves a Reasoning score of 89.7, outperforming GPT-4-turbo and Claude-3.5-sonnet, with a substantial 15.9-point gain over baselines. Dr.SoW focused on chat-hard scores 69.7 in ChatHard, reflecting improved reward robustness in challenging dialog contexts.
Dr.SoW uses an oracle (idealized routing) to establish a performance upper-bound with dynamic routing. Under ideal conditions, it achieves an overall score of 84.9, balancing safety, reasoning, and conversational robustness. In practice, adaptive Dr.SoW employs a router (a zero-shot LLM) to automate domain assignment. Notably, the router uses the vanilla density ratio for the general chat domain, as it performs best in Chat, which is the most frequent scenario in real-world annotation settings.
Overall, Dr.SoW outperforms standard density ratio baselines by as much as 5.4 points, showing the advantages of adaptively customized reward functions. Generative reward using the same strong model with an identical instruction set performs near random chance. In contrast, Dr.SoW that contrasts the strong model versus a weaker model achieves 82.6 overall. The performance is comparable to LLM-as-a-judge reward from GPT-4-turbo and Claude-3.5-sonnet, and surpasses the best in-class Mistral-7B classifier reward.
@@ 4.3 Alignment with Density Ratio Annotated Data
Previous experiments indicated that Dr.SoW delivers a strong reward signal, achieving high scores on standard reward benchmarks. Here, we preference-tune LLMs using data annotated by Dr.SoW, enabling direct comparisons between Dr.SoW and SoTA reward functions in their effectiveness for preference alignment.
@@
We initialize with Meta-Llama-3-8B-Instruct and preference-tune it using SimPO (Meng et al., 2024) with data annotated by Dr.SoW, along with other reward functions (see Appendix A.1 for details). Details about the SimPO algorithm and our training setup are available at Appendix A.2. Our evaluation methods include AlpacaEval2.0, ArenaHard, and MT-Bench (details in Appendix B).
@@
Dr.SoW (safety) 30.0 44.7 2850 39.4 777 8.0
Dr.SoW (code/math) 36.0 33.1 1853 30.4 545 8.2
Dr.SoW (adaptive) 40.7 46.1 2229 37.4 643 8.0
Table 2: Alignment performance after SimPO training on the Llama-3-Instruct (8B) model. Reward function is used to annotate the online preference dataset, obtained through Best-of-32 sampling. The first row is the performance of the starting model Llama-3-Instruct (8B) model. The second row is the alignment performance of aligning using a SoTA trained reward function. DPO model indicated is NousResearch/Nous-Hermes-2-Mistral-7B-DPO; SFT model is teknium/OpenHermes-2.5-Mistral-7B; Base model is mistralai/Mistral-7B-v0.1. Dr.SoW applies domain-specific guidance (e.g., safety or code/math) to the vanilla density ratio reward. Adaptive indicates using a routing system to assign domain-related instruction set for each example.
@@
We focus on two model pairs in the Dr.SoW reward formulation: (i) SFT vs. Base, and (ii) DPO vs. SFT. The first model pair (SFT vs. Base) is chosen because neither model has undergone preference tuning, allowing us to test whether a preference reward can be derived based purely on the overall capability improvement after SFT training. The second model pair (DPO vs. SFT) is selected for its reward performance, as shown in Table 1. For the prompt-guided reward function, we experiment with various instruction types: no instructions, safety domain instructions, math/coding domain instructions, and adaptive instructions tailored to the domain of each input prompt.
Results
As shown in Table 2, Llama-3-instruct preference fine-tuned using data annotated by the DPO-over-SFT density ratio achieve strong performance, with 39.9 on AlpaceEval 2 and 34.6 on ArenaHard. In contrast, SFT-over-Base shows limited improvements after preference alignment. Narrow gap in their human-aligned level results in noisy reward signal that fails to annotate preference data effectively. This demonstrates again that the effectiveness of reward function in (2) depends on a significant gap in human-value alignment between the numerator and denominator models.
Table 2 shows that reward functions customized for specific domain can not be applied universally to all examples, doing so would result in suboptimal performance, as in “safety” and “code/math” Dr.SoW results. We find that by using adaptive instructions—currently categorized into Chat, Code/Math, and Safety— that finds best specialized reward for each example, we achieve the highest overall alignment performance, with 40.7 on AlpacaEval 2 and 37.4 on ArenaHard, competitive against SoTA reward from ArmoRM. Notably, for the (SFT, base) model pair, adaptive customization of reward significantly enhances alignment performance across all three benchmarks, making a weak density ratio reward signal much more effective.
5 Related Works
Preference tuning
Many preference tuning algorithms have been proposed to align LLMs with human preferences and values (Melnyk et al., 2024; Pang et al., 2024; Ethayarajh et al., 2024; Wu et al., 2024; Hong et al., 2024; Yuan et al., 2023). The most well-known one is the proximal policy optimization (PPO; Schulman et al., 2017), an online RL algorithm that optimizes policy to maximize the KL-constrained reward expectation of an external reward model. Direct preference optimization (DPO; Rafailov et al., 2023a) leverages DPO implicit reward – parameterized as density ratio between policy model and a reference model—to circumvent the need of external reward function. It simultaneously optimizes the implicit reward and policy model by training on pairwise preference data. More recently, SimPO (Meng et al., 2024) directly optimizes the average log-likelihood margin between winning and losing sequences, eliminating the need for a reference model.
Density ratio reward functions
Density ratio as reward function is popularized by implicit DPO reward  (Rafailov et al., 2023a).  Chen et al. (2024) uses implicit DPO reward to bootstrap an LLM through iterative DPO training. Zhong et al. (2024) trains a DPO model and uses the density ratio to derive a token-level characterization for response quality, and uses it as a reward signal in PPO training. Yang et al. (2024b) uses the density ratio between DPO vs SFT model as quality filter. Though one study Lin et al. (2024) finds that implicit DPO reward struggles to generalize on OOD examples compared with just training a classifier using (BradleyTerry; Bradley & Terry, 1952) objective. This work extends the density ratio reward formulation to broader spectrum of models, and provides guidance for finding stronger reward signal than implicit DPO reward.
Discriminative & generative rewards
Trained classifiers and generative rewards are the mainstream method for preference data annotation. They top leaderboards such as RewardBench  (Lambert et al., 2024) and are widely used to preference align well-known models (Ouyang et al., 2022; Touvron et al., 2023; Adler et al., 2024; Yang et al., 2024a; Cui et al., 2023). High quality and popular preference datasets are often annotated using powerful proprietary models as-a-judge, either in the forms of scalar score or textual assessment and critiques (Cui et al., 2023). Then, one can use the data to finetune a generative judge (Wang et al., 2024b; Zhang et al., 2024; Wang et al., 2024a; Kim et al., 2024) or to train a sequence classifier (Adler et al., 2024; Dong et al., 2024; Liu & Zeng, 2024). Dr.SoW provides a data-free and training-free alternative for reward modeling and preference annotation.
Weak-to-strong generalization
Prior works have explored the idea of contrasting a weak and a strong model to obtain better performance than the strong model. Contrastive decoding (CD), for instance, enhances LLM generation quality by searching for sequences that maximizes the likelihood difference between an expert model and an amateur model. O’Brien & Lewis (2023) shows CD consistently improves reasoning tasks.  Li et al. (2022) shows improved generation quality in wikipedia, news and story domains.  Chuang et al. (2023) shows improvement in LLM facutuality by contrasting the differences between logits in later layers and earlier layers. ExPo (Zheng et al., 2024) uses the delta between an aligned model and pre-aligned model to extrapolate a better aligned models through weight merging. Dr.SoW similarly contrasts strong-over-weak models, and uses the delta to align small models to near GPT-4 level performance on ArenaHard (Figure 13).
6 Conclusion and Future Work
We introduce Dr.SoW, a cost-effective and accessible approach that uses off-the-shelf LLMs for preference data annotation. It reduces the need for costly human labeling or proprietary models to achieve a high-performance reward function. At the core of Dr.SoW is the Strong-over-Weak hypothesis, which we rigorously validate through extensive experiments. This insight offers a design guideline for practitioners seeking LLM-based preference annotation.
Domain-specific customization further enhances the density ratio reward, particularly in targeted areas such as safety and reasoning. And this is achieved without requiring additional data or fine-tuning. We offer an automated pipeline to adaptively combine domain-expert reward functions for tailored preference annotation. This approach shows strong performance on reward benchmarks, and its annotated data pushes an 8B model to GPT-4 level performance on ArenaHard (Figure 13). This result is competitive with state-of-the-art (SoTA) reward classifiers while avoids the data and compute overheads of actually training reward functions, highlighting Dr.SoW as both cost-effective and highly effective.
Recently, density ratio based reward functions have demonstrated state-of-the-art performance as Math Process-Reward Models (PRMs) (Yuan et al., 2024), as it provides token-level value estimates. Exploring the use of Dr.SoW for process-level presents a promising future direction, particularly for inference-time scaling use-cases.
Appendix A Experimental Details
A.1 Preference Data Annotation
A.2 Training Details